OpenStudy (anonymous):

how do you find the value of 24xyz

OpenStudy (zzr0ck3r):

we need more info

OpenStudy (anonymous):

so the question was 2x=5 3y=4 4z=3 what is the value of 24xyz

OpenStudy (zzr0ck3r):

solve for x solve for y solve for z then tell me what x=? y=? z=? like this 2x=5 so x=5/2

OpenStudy (zzr0ck3r):

so do the same thing for y, and z.

jimthompson5910 (jim_thompson5910):

you could also multiply the left sides together then multiply the right sides together

OpenStudy (zzr0ck3r):

that's prob more clever way yes;)

OpenStudy (zzr0ck3r):

@RebaLynn do you see what he means?

OpenStudy (jhannybean):

(2x)(3y)(4z) = (5)(4)(3) ?

OpenStudy (zzr0ck3r):

2x=5 3y=4 4z=3 multiply everything on the left hand sides 2x*3y*4z = 24xyz and since we multiplies everything on the left hand sides, this will equal what we get when we multiply the right hand sides

OpenStudy (anonymous):

so I just divide 2x\5 and get the value of x

OpenStudy (zzr0ck3r):

that is one way method one 2x=5 3y=4 4z=3 so x=5/2, y=4/3, z=3/4 so 24xyz=24(5/2)*(4/3)*(3/4) = (24*5*4*3)/(2*3*4) = 12*5 = 60 method two multiply the left hand sides and the right hand sides together (2x)(3y)(4z) = (5)(4)(3) 24xyz=60

OpenStudy (jhannybean):

So just find the value of xyz. and then plug it in to find 24(xyz)

OpenStudy (anonymous):

if the average of 5 numbers is 10 and the fitfth number Is 6 then what is the average of the first 4 numbers

OpenStudy (anonymous):

average is the sum of all the number DIVIDED by the amount of numbers...so if the average of 5 numbers is 10, multiply 10 by 5, which would give you that the SUM of the 5 numbers is 50 now if the 5th number is 6, you SUBTRACT 6 from 50 to get that the first 4 numbers add up to 44..now just divide by 4 to get the average of the first FOUR numbers...so the average of the first four numbers would be 11 :)